OpenStudy (anonymous):

solve 2x^2=10 and find x intercepts of f(x)=2x^2-10

OpenStudy (anonymous):

Divide each side by 2 2x^2/2 = 10/2 x^2 = 5 square root (x^2) = square root (5) x = +/- (square root) 5

OpenStudy (anonymous):

ok and what about the x intercepts?

OpenStudy (anonymous):

You create a table and insert for x X| 1 | 2 | 3 | 4 | 5 | Y| -6| 6| etc.....

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

here it shows a graph im going to put it up maybe you can help

OpenStudy (anonymous):

2x^2=10 ===> x^2 = 10/2 =5 ===> x= sqrt(5) or x=-sqrt(5)

OpenStudy (anonymous):

btw sqrt(5) = 2.24

OpenStudy (anonymous):

ok and can you give me the x intercepts?

OpenStudy (anonymous):

what we solved are the x-intercept

OpenStudy (anonymous):

so its (0,2.24)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

solve 2x^2=10 and find x intercepts of f(x)=2x^2-10

OpenStudy (anonymous):

swich the coordinates ( sqrt(5),0) =(2.24,0) , (-sqrt(5),0)=(-2.24,0)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

hope this helps.

OpenStudy (anonymous):

Thanks